40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5233     Entry time: Sun Aug 14 20:04:40 2011
Author: Keiko, Anamaria, Jenne, and Kiwamu 
Type: Summary 
Category: Locking 
Subject: central part ifo locking plan 
GOAL : To lock the central part of ifo

Here is the plan:

Mon - assemble all the cables from PDs and mixers, and check the CDS channels. Prepare the beamsplitters.

Tue - The current paths to REFL11 and REFL55 will be modified to the four paths to REFL11, 33, 55, 165. And the PDs will be placed.
Wed, Thu - during waiting for the ifo available with vacuum, help aligning the POP, POX, POY. In parallel, a simulation to find the PRC length SRC 
length tolerance will be proceeded.

Fri - When the ifo becomes available with vacuum, the sensing signals by 3-f scheme will be obtained with proper demodulation phases.

Sat - Try to lock the central part of the ifo with the new 3-f signals.
ELOG V3.1.3-